Give the necessary theory and method for determining the surface tension of water by capillary rise method.  


1
Expert's answer
2021-03-10T09:35:44-0500

According to theory, the capillary rise is


"h
=\frac{
2
σ
\text{cos}
θ}{
ρ
g
r},
\\\space\\
\sigma=\frac{h\rho gr}{2\text{cos}\theta}."

So, we need to measure the radius of the capillary tube r, know the density of liquid, and measure the angle theta: